# Deepfake-Eval-2024: A Multi-Modal In-the-Wild Benchmark of Deepfakes Circulated in 2024

### v2

@@ 4.2 Finetuned Model Performance
Most models improve considerably when finetuned on a subset of Deepfake-Eval-2024. AUC improves by an average of 57.6% for video, 80.6% for audio and 15.6% for images (Table 6, Supp. Table S5). However, the degree of improvement varies across models, suggesting that some model architectures may be less suited to adapt to the challenges of real-world deepfake detection. For example, the simple single-layer UFD model learns to predict all data as fake after finetuning, and Styleflow and DistilDIRE also show limited improvement in AUC. However, this limited improvement could also be attributed to the relatively small finetuning set size. Although in most cases performance improves after finetuning, there is still significant room for improvement, with the peak accuracy reaching 0.75 for videos, 0.86 for audio, and 0.69 for images, which is lower than the 90% lower bound of human deepfake analyst accuracy that we estimate from inter-labeler agreement (Table 4). These results suggest that in addition to more representative training datasets, new model paradigms may be needed for robust and reliable deepfake detection.
@@ 4.3 Commercial Model Performance
Top commercial models exceed the performance of open-source models. Top commercial models considerably outperform off-the-shelf open-source models and finetuned image models, and perform slightly better than finetuned audio and video models. No commercial models that we evaluated had an accuracy of 90% or above, suggesting that commercial models still need improvement to reach accuracy of human deepfake forensic analysts. In addition, we note that open-source models finetuned on a subset of Deepfake-Eval-2024 approach the accuracy of commercial models (finetuned video model GenConViT has an accuracy of 75%, and finetuned audio model P3 from [38] has an accuracy of 86%). This suggests that the competitive advantage of these commercial models may be derived primarily from training dataset curation.
4.4 Cross Modality Trends
In both open-source and commercial models, performance across modalities varies, with audio detectors consistently performing with higher accuracy than image and video detectors. This suggests that audio deepfakes may currently be more algorithmically distinguishable than video or image deepfakes, possibly due to the more constrained nature of audio manipulation techniques. However, the strong performance of audio models may be confounded by the fact that commercial audio model detectors were used during labeling to help distinguish between real voice impersonators and AI-generated audio, thus potentially biasing the audio dataset towards a greater quantity of audio that is already easily detected by existing audio models.
@@ 5 Error Analysis
To further investigate deepfake detection model failures, we identify media traits associated with errors made by open-source models on Deepfake-Eval-2024. Our error analysis identifies media characteristics that current models fail on, exhibiting traits present in Deepfake-Eval-2024 that were likely missing from detection model training and testing datasets.
Error analysis methodology. We perform manual error analysis on the entire finetuning test set of videos and images. Due to the length of the Deepfake-Eval-2024 audio test set we were unable to manually analyze its entirety, and instead manually evaluate a class-balanced random sample of 10%, consisting of 2000 four-second audio clips. For each modality we identify media traits that are associated with a statistically significant decrease in accuracy as measured by chi-squared tests for each model separately, using significance threshold $p<0.05$. We exclude models that predict all data as belonging to a single class from this error analysis (off-the-shelf and finetuned UFD [35] and off-the-shelf P3[38]).
@@
Models perform worse on audio with non-English languages, silences, and background noise, specifically music. We focus exclusively on the errors of finetuned audio models, because due to the low accuracy of off-the-shelf audio models (Table S5), floor effects make off-the-shelf performance indistinguishably poor across traits. In finetuned models, non-English audio has an average accuracy that was 7.21% lower than English audio. Because Deepfake-Eval-2024 is an in-the-wild dataset, some parts of the included audio files are silent. Models have 35.39% worse accuracy on audio clips that are silent. This is expected behavior, as audio without speakers is out of distribution for models targeting deepfake audio. We also note that models perform worse on audio with background noises (accuracy decrease of 7.66%). Music in the background is associated with a drop in accuracy of 17.94%, and a large increase of 26.12% in false negative rate. Adding background music is a common technique in deepfake generation, and our results suggest that current models often fail to identify fakes when the fakes have music added. The inability of models to accurately predict audio with music is a major vulnerability in existing audio deepfake detection models.
@@
Models have lower accuracy on images with text overlays. For images, we identify characteristics such as images containing crowds, depicted skin color, and the presence of text overlays. We did not find any statistically significant differences in performance associated with these categories, but we do observe a notable decrease in accuracy when images have text: accuracy of finetuned models decreases by 9% and the F1-score decreases by 10.5% on average. This indicates a distributional mismatch with existing training datasets, which do not include images with text overlays. We hypothesize that the lack of statistically significant differences in the image models is due to floor effects, as the average accuracy of off-the-shelf and fine-tuned audio models was low (Tables S5, S5).
@@
Many errors are not attributable to human-identifiable characteristics. The identified error-associated media characteristics do not encompass all errors in the dataset. Only a fraction of the entire dataset has each of these traits, and thus only a fraction of the errors have these traits as well (e.g. an average of 33% of audio errors have music in the background, an average 5% of the image errors had text overlays), and there are model errors on media with none of the error-associated traits. As such, we conclude that many errors are not associated with human-identifiable characteristics but are instead caused by failures in model signal interpretation or other invisible artifacts. Discovering the signal patterns responsible for detection errors is an important area of future research.
6 Discussion
Conclusion. We present Deepfake-Eval-2024, an in-the-wild deepfake dataset that captures the challenging real-world threat of contemporary deepfakes through a collection of diverse and recent data gathered from social media and TrueMedia.org users around the world. State-of-the-art models perform far worse on Deepfake-Eval-2024 than on previous benchmarks, suggesting that past popular deepfake datasets are not representative of the contemporary challenge of deepfake detection, and that off-the-shelf open-source deepfake detection models are not capable of accurately detecting real-world deepfakes.
Future work. While Deepfake-Eval-2024 is the most comprehensive and diverse collection of real-world deepfakes available today, the rapid evolution of generative AI necessitates the development of robust monitoring systems to track emerging deepfakes and regularly update datasets. We also acknowledge that creating manually labeled in-the-wild deepfake datasets is costly and susceptible to labeling errors. As such, we recommend that future synthetic datasets should strive to be more representative of real-world data, and contain characteristics that our study reveals to be associated with model errors. Our error analysis illustrates that we also need automated methods to identify model errors before they can be exploited. Ultimately, we believe the release of Deepfake-Eval-2024 as a benchmark for contemporary real-world deepfake detection will catalyze the development of robust models that can effectively address the evolving threat of modern deepfakes.
@@ A.2 Dataset Supplementary Figures
Figure S1: Origins of data in Deepfake-Eval-2024 combined for all modalities. Direct upload indicates that the media was uploaded directly to TrueMedia.org by a user, instead of the user providing a link to a social media website.
@@
A.3 Results Supplementary Figures
@@
NPR 0.73 0.69 0.74 0.78 0.76 0.46 0.22 -
@@ Appendix B Appendix: Labeling Criteria
We present the labeling criteria for all modalities. The complementary examples mentioned in this section can be found at this GitHub link.
@@ B.3 Audio Codebook
Music, silence, and sound effects were labeled as real unless there was other evidence of AI manipulation. Audio-Only: if 2 or more models say >80% PLUS there’s some additional reason to believe it’s fake (ie. the audio quality sounds synthetic) Example 11 Voice is off camera and unable to locate original source

### v4

@@ 4.2 Finetuned Model Performance
Models improve when finetuned on a subset of Deepfake-Eval-2024. AUC improves by an average of 57.6% for video, 80.6% for audio and 4.5% for images (Table 4, Supp. Table S8). However, the degree of improvement varies across models, suggesting that some model architectures may be less suited to adapt to the challenges of real-world deepfake detection. For example, the simple single-layer UFD model learns to predict all data as fake after finetuning, and video model Styleflow also shows limited improvement in AUC. The limited improvement in image models is likely attributable to the relatively small amount of image finetuning data, which was insufficient to shift the highly paramaterized models of NPR and DistilDIRE towards the distribution of in-the-wild deepfakes. Although performance improves after finetuning, there is still significant room for further improvement, with the peak accuracy reaching 0.75 for videos, 0.86 for audio, and 0.63 for images, which is below the 90% lower bound estimate of human deepfake analyst accuracy (Section 3.3). These results suggest that in addition to more representative training datasets, new model paradigms may be needed for robust and reliable deepfake detection.
@@ 4.3 Commercial Model Performance
Top commercial models exceed the performance of open-source models. Top commercial models considerably outperform off-the-shelf open-source models and finetuned image models, and perform slightly better than finetuned audio and video models. Commercial audio model performance is particularly strong, which is likely attributable to the limitations of the audio labeling approach as described in 3.3. No commercial models that we evaluated had an accuracy of 90% or above, suggesting that commercial models still need improvement to reach the accuracy of human deepfake forensic analysts. In addition, we note that open-source models finetuned on a subset of Deepfake-Eval-2024 approach the accuracy of commercial models (finetuned video model GenConViT has an accuracy of 75%, and finetuned audio model P3 from [42] has an accuracy of 86%). This suggests that the competitive advantage of these commercial models may be derived primarily from training dataset curation.
@@ 5 Error Analysis
Error analysis methodology. To further investigate detection model failures, we identify media traits associated with errors. We perform manual error analysis on the entire finetuning test set of videos and images. Due to the length of the Deepfake-Eval-2024 audio test set, we were unable to evaluate its entirety and instead we manually evaluate a class-balanced random sample of 10%, consisting of 2000 four-second audio clips. For each modality we identify media traits that are associated with a statistically significant decrease in accuracy as measured by chi-squared tests for each model separately, using significance threshold $p<0.05$. We exclude models that predict all data as belonging to a single class from this error analysis (off-the-shelf and finetuned UFD [39] and off-the-shelf P3[42]).
@@
Models perform worse on audio with non-English languages, silences, and background noise, specifically music. We focus exclusively on the errors of finetuned audio models because floor effects associated with low performance (Table S8) make off-the-shelf performance indistinguishably poor across traits. In finetuned models, non-English audio has an average accuracy that was 7.21% lower than English audio. Because Deepfake-Eval-2024 is an in-the-wild dataset, some parts of the included audio files are silent. Models have 35.39% worse accuracy on audio clips that are silent. This is expected behavior, as audio without speakers is out of distribution for models targeting deepfake audio. We also note that models perform worse on audio with background noises (accuracy decrease of 7.66%). Music in the background is associated with a drop in accuracy of 17.94%, and a large increase of 26.12% in false negative rate. Adding background music is a common technique in deepfake generation, and our results suggest that current models often fail to identify fakes when the fakes have music added. The inability of models to accurately predict audio with music is a major vulnerability in existing audio deepfake detection models.
@@
Models have lower accuracy on images with text overlays. We identify image categories such as depicted crowds, skin color, and the presence of text overlays. We did not find any statistically significant differences in performance associated with these categories, but we do observe a decrease in accuracy when images have text: accuracy of finetuned models decreases by 9% and the F1-score decreases by 10.5% on average. This indicates a distributional mismatch with existing training datasets, which do not include images with text overlays. We hypothesize that the lack of statistically significant differences in the image models is due to floor effects, as the average accuracy of off-the-shelf and fine-tuned audio models was low (Tables S8, S8).
@@
Many errors are not attributable to human-identifiable characteristics. The identified error-associated media characteristics do not encompass all errors in the dataset (e.g. an average of 33% of audio errors have music in the background, and an average 5% of the image errors had text overlays). The errors not associated with media traits are caused by other failures in model signal interpretation. Developing methods to identify non-visible error patterns is an important area of future research.
6 Discussion & Limitations
In-the-wild deepfake data is crucial for evaluating detection models against real-world threats. However, we acknowledge that curating and manually labeling in-the-wild data is costly and susceptible to human error, resulting in a dataset size appropriate for evaluation but insufficient for wide-scale training. As such, there is still a need for synthetic deepfake datasets, and we recommend that these should strive to be more representative of real-world data and contain the media characteristics that our study reveals to be associated with model errors.
While Deepfake-Eval-2024 is the most comprehensive and diverse collection of real-world deepfakes available today, the rapid evolution of generative AI means that datasets can quickly become outdated, and more work is needed to develop systems that track emerging deepfakes and regularly update datasets. There is also a risk of adversarial actors using Deepfake-Eval-2024 to develop new deepfake generation techniques that evade detectors. To address this risk we make the dataset available at https://huggingface.co/datasets/nuriachandra/Deepfake-Eval-2024 with a CC-BY-SA-4.0 license, but gate access to individuals who are verifiably working on deepfake detection or at research institutions (See Appendix A.5).
Ultimately, we believe the release of Deepfake-Eval-2024 as a benchmark for contemporary real-world deepfake detection will catalyze the development of robust models that can effectively address the evolving threat of modern deepfakes.
@@ A.2 Dataset Supplementary Figures
Table S4: Deepfake-Eval-2024 Video Summary Statistics
Table S5: Deepfake-Eval-2024 Audio Summary Statistics
Real 36.6 1,110 124.80 44.83
Fake 19.9 710 101.51 44.40
All 56.5 1,820 115.46 44.66
Table S6: Deepfake-Eval-2024 Image Summary Statistics
Category Count Mode Resolution
Fake 1,208 1,200$\times$1,200
Real 767 1,024$\times$1,024
All 1,975 1,024$\times$1,024
Figure S1: Origins of data in Deepfake-Eval-2024 separated by modality. In total, media was shared from 88 different web-domain names. Direct upload indicates that the media was uploaded directly to TrueMedia.org by a user, instead of the user providing a link to a social media website.
@@
A.3 Experimental Compute
We evaluated and finetuned all models on a single AWS A10 GPU on an g5.2xlarge instance or a single GCP L4 GPU on a g2-standard-8 instance. We found no difference between evaluating and finetuning models on the AWS or GCP instance.
A.4 Results Supplementary Figures
@@
NPR 0.55 0.58 0.61 0.81 0.70 0.76 0.19 -
A.5 Dataset Access
Deepfakes pose an established threat to society, and there is a potential for Deepfake-Eval-2024 to be used with malicious intentions to create deepfake generation technologies that are more realistic and that evade existing detectors. As such, we gate access to this dataset through Huggingface to individuals verifiably at research institutions or doing work related to deepfake detection. This is done via a Huggingface access request interface where we request an institutional / company email address. There is also a space for individuals to provide additional evidence of work related to deepfake detection, which is used for verification in the case that the email address provided is a) not connected to an institution or b) is not in the name of the requesting individual. We intend to maintain this system of access gating until May 2027, at which point generative AI is likely to have advanced far beyond the deepfakes represented in Deepfake-Eval-2024, and thus there will be a much lower risk of making Deepfake-Eval-2024 available to all individuals. Starting in May 2027, we will modify the access process so that any individual who agrees to the Terms of Use can get access to the dataset.
@@ Appendix B Appendix: Labeling Criteria
We present the labeling criteria for all modalities. The complementary examples mentioned in this section can be found in Supplementary .zip file.
@@ B.3 Audio Codebook
Audio without speech such as music, silence, and sound effects were labeled as real unless there was other evidence of AI manipulation. Audio-Only: if 2 or more models say $>80$% likelihood of fake PLUS there’s some additional reason to believe it’s fake (ie. the audio quality sounds synthetic, or sociocultural implausability) Example 11 Voice is off camera and unable to locate original source